# Consistency of Neural Causal Partial Identification

### v1

@@ B.1 Proof of Theorem 1
$(\gamma_{n+1})_{\#}\lambda(C_{O_{n+1}(i)})=\mathbb{P}(C_{O_{n+1}(i)}),(\gamma_{n+1})_{\#}\lambda(\partial C_{O_{n+1}(i)})=0,\forall i.$ (18)
The idea is to construct a piecewise linear curve going through all cubes in $\mathcal{C}_{n+1}$ exactly once and modify its speed. Take $v_{0}\in\text{Center}(C_{O_{n+1}(1)}),v_{i}\in\text{Center}(C_{O_{n+1}(i)}\cap C_{O_{n+1}(i+1)}),v_{2^{d(n+1)}}\in\text{Center}(C_{O_{n+1}(2^{d(n+1)})})$, where $\text{Center}(\prod_{i=1}^{k}[a_{i},b_{i}])=((a_{i}+b_{i})/2)_{i=1,\cdots,k}$ is the center of a cube. $v_{i}$ is well-defined since $C_{O_{n+1}(i)}\cap C_{O_{n+1}(i+1)}$ are cubes of dimension $d-1$ by Lemma 1. One possible choice of $\gamma_{n+1}(t)$ is
$\gamma_{n+1}(t)=v_{0}+\sum_{i=1}^{2^{(n+1)d}}\frac{1}{p_{i}}(v_{i}-v_{i-1})\left(\left(t-\sum_{k=1}^{i-1}p_{k}\right)^{+}-\left(t-\sum_{k=1}^{i}p_{k}\right)^{+}\right).$ (19)
@@
Finally, for any $k_{1}\geqslant k_{2}\geqslant n,t\in[0,1]$, by property 2 in step 1, (17) and (19), $\gamma_{k_{1}}(t),\gamma_{k_{2}}(t)$ are in one cube $C_{i_{1},i_{2}\cdots,i_{n}}$. and $\text{diam}(C_{i_{1},i_{2}\cdots,i_{n}})\leqslant\left(\frac{2}{3}\right)^{n}\sqrt{d}$. Thus, $|\gamma_{k_{1}}(t)-\gamma_{k_{2}}(t)|\leqslant\left(\frac{2}{3}\right)^{n}\sqrt{d}$ which implies $\{\gamma_{n}(t)\}$ converges uniformly to one continuous function $\gamma(t)$.
@@
$=\sum_{i_{n+1},\cdots,i_{k}}\mathbb{P}(C_{i_{1},i_{2}\cdots,i_{k}})=\mathbb{P}(C_{i_{1},i_{2}\cdots,i_{n}}),$ (20)
@@
where we use (20) and the fact that $(\gamma_{k})_{\#}\lambda$ vanishes on $\partial C_{i_{1},i_{2}\cdots,i_{k}}$ in the first equality and $\text{diam}(C_{i_{1},i_{2}\cdots,i_{k_{1}}})\leqslant\left(\frac{2}{3}\right)^{k_{1}}\sqrt{d}$ in the last inequality. Let $k\rightarrow\infty$, by Portmanteau Theorem, we get
@@
and we conclude that $\gamma_{\#}\lambda(\partial C_{i_{1},i_{2}\cdots,i_{n}})=0$. By Portmanteau Theorem, let $k\rightarrow\infty$ in (20), we obtain $\gamma_{\#}\lambda(C_{i_{1},i_{2}\cdots,i_{n}})=\mathbb{P}(C_{i_{1},i_{2}\cdots,i_{n}})$. Let $\mathcal{C}^{*}=\bigcup_{i=1}^{\infty}\mathcal{C}_{i}$. Notice that for any open set $U\in[0,1]^{d}$, $U=\bigcup_{C\subset U,C\in\mathcal{C}^{*}}C$, which means the $\sigma$-algebra generated by $\mathcal{C}^{*}$ contains all Borel sets. Besides, $\mathcal{C}^{*}$ is a $\pi$-system. Hence, $\lambda\left(\gamma^{-1}(U)\right)=\mathbb{P}(U)$ for all Lebesgue measurable set $U$.
Step 4: Holder Continuity of $\gamma$. We verify the condition of [46, Theorem 1]. $\mathcal{I}_{n}$ and $\mathcal{C}_{n}$ in the above construction correspond to developments $\alpha_{k},\beta_{k}$ in [46, Theorem 1]. Define $f_{k}(I_{O_{k}(i)})=C_{O_{k}(i)},\ i=1,\cdots,2^{kd}$ as a map from $\mathcal{I}_{n}$ to $\mathcal{C}_{n}$.
@@
    If $I_{1},I_{2}\in\mathcal{I}_{k}$ and $I_{1}\cap I_{2}\neq\emptyset$, $I_{1}$, $I_{2}$ are adjacent. By Lemma 1, $f_{k}(I_{1}),f_{k}(I_{2})$ share a $(d-1)$-dimensional boundary. Therefore, $f_{k}(I_{1})\cap f_{k}(I_{2})\neq\emptyset$.
@@
    Let $A=\prod_{k=1}^{\infty}\left(1+2^{-k}\right)$. We have for any $C\in\mathcal{C}_{n}$, by 5
@@
Proof of Theorem 3.
The proof is similar to the proof of Theorem 2. The difference is that in the first part, we use a deep ReLU network to approximate the Holder continuous curve $\gamma$ instead of wide ReLU network.
Let $\mu_{i}=\mathbb{P}_{C_{i}}/\mathbb{P}(C_{i})$. By 4 and 5, for each connected component $C_{i}$ of the support, there exists Lipschitz maps $g^{i}_{2}\in\mathcal{H}([0,1]^{d^{C}_{i}},C_{i})\cap\mathcal{F}_{L}([0,1]^{d^{C}_{i}},C_{i})$ such that
@@
for any measurable set $B\subset[0,1]^{d_{i}^{C}}$. By Proposition 1, there exist a $1/d_{i}^{C}$-Holder continuous curve $\gamma_{i}:[0,1]\rightarrow[0,1]^{d_{i}^{C}}$ such that $(\gamma_{i})_{\#}\lambda=(g^{i}_{2})^{-1}_{\#}\mu_{i}$.
By [52, Theorem 2] (take $\omega(x)=C\|x\|^{1/d}$ where $C$ is the Holder continuity constant factor), there exists deep ReLu network $\hat{g}_{1}^{i,j}$ of width $12$ and depth $L_{1}$ such that
$\|\hat{g}_{1}^{i,j}-(\gamma_{i})_{j}\|_{\infty}\leqslant O\left(L_{1}^{-2/d_{i}^{C}}\right),$
where $(\gamma_{i})_{j}$ is the $j$-coordinate of $\gamma_{i}$. Let $\hat{g}_{1}^{i}(x)=\left(\hat{g}_{1}^{i,1}(x),\cdots,\hat{g}_{1}^{i,d}(x)\right)$, we get
$\|\hat{g}_{1}^{i}-\gamma_{i}\|_{\infty}\leqslant O\left(L_{1}^{-2/d_{i}^{C}}\right).$
Thus, the width of $\hat{g}_{1}^{i}$ is $\Theta\left(d_{i}^{C}\right)$. By Lemma 3
$W\left(\left(g_{2}^{i}\right)^{-1}_{\#}\mu_{i},P(\hat{g}_{1}^{i}(U_{i}))\right)$ $=W\left((\gamma_{i})_{\#}U_{i},P(\hat{g}_{1}^{i}(U_{i}))\right)\leqslant\|\hat{g}_{1}^{i}-\gamma_{i}\|_{\infty}\leqslant O\left(L_{1}^{-2/d_{i}^{C}}\right).$
The rest are the same as proof of Theorem 2. ∎
Proof of Corollary 1.
@@
$V_{i}=\begin{cases}f_{i}\left(\text{Pa}(V_{i}),\boldsymbol{U}_{V_{i}}\right),&V_{i}\ \text{is continuous},\\ \arg\max_{k\in[n_{i}]}\left\{g_{k}^{V_{i}}+\log\left(f_{i}\left(\text{Pa}(V_{i}),\boldsymbol{U}_{V_{i}}\right)\right)_{k}\right\},&V_{i}\ \text{is categorical},\|f_{i}\|_{1}=1\end{cases},\quad f_{i}\in\mathcal{F}_{L}.$
@@
$\hat{f}_{i}=\begin{cases}\arg\min_{\hat{f}_{i}\in\mathcal{NN}_{d_{i}^{\text{in}},d_{i}^{\text{out}}}(d_{i}^{\text{out}}(2d_{i}^{\text{in}}+10),L_{0})}\|\hat{f}_{i}-f_{i}\|_{\infty},&V_{i}\ \text{continuous},\\ \max\{0,\arg\min_{\hat{f}_{i}\in\mathcal{NN}_{d_{i}^{\text{in}},d_{i}^{\text{out}}}(d_{i}^{\text{out}}(2d_{i}^{\text{in}}+10),L_{0})}\|\hat{f}_{i}-f_{i}\|_{\infty}\},&V_{i}\ \text{categorical}.\end{cases}$
@@
We truncate the neural network at 00 for categorical variables because the propensity functions are required to be non-negative. This truncate operation will not influence the approximation error since $f_{i}$ are non-negative. According to 3, $f_{i}$ are Lipschitz continuous. By [52, Theorem 2], we have, if $V_{i}$ is continuous, $\|\hat{f}_{i}-f_{i}\|_{\infty}\leqslant O(L_{0}^{-1/d_{i}^{\text{in}}})\leqslant O(L_{0}^{-1/d_{\max}^{\text{in}}})$.
If $V_{i}$ is categorical, let $S=\|\hat{f}\left(\text{pa}(v_{k}),\boldsymbol{u}_{V_{k}}\right)-f\left(\text{pa}(v_{k}),\boldsymbol{u}_{V_{k}}\right)\|_{\infty}$, $\boldsymbol{p}={f}\left(\text{pa}(v_{k}),\boldsymbol{u}_{V_{k}}\right),\hat{\boldsymbol{p}}=\hat{f}\left(\text{pa}(v_{k}),\boldsymbol{u}_{V_{k}}\right)/\|\hat{f}\|_{1}$. If $S>1/(2K)$, $\|\boldsymbol{p}-\hat{\boldsymbol{p}}\|_{\infty}\leqslant 2KS$. If $S\leqslant 1/(2K)$,
$\|\boldsymbol{p}-\hat{\boldsymbol{p}}\|_{\infty}$ $\leqslant\frac{1}{\|\hat{f}\|_{1}}\|\hat{f}\left(\text{pa}(v_{k}),\boldsymbol{u}_{V_{k}}\right)-f\left(\text{pa}(v_{k}),\boldsymbol{u}_{V_{k}}\right)\|_{\infty}+\frac{1}{\|\hat{f}\|_{1}}|\|\hat{f}\|_{1}-1|\|f\|_{\infty}$
$\leqslant\frac{(n_{i}+1)}{\|\hat{f}\|_{1}}S\leqslant\frac{(n_{i}+1)}{1-n_{i}S}S\leqslant 2(K+1)S$
where we use 3 that $n_{i}\leqslant K$ and $S\leqslant 1/(2K)$. Thus,
$\|f_{i}-\hat{f}_{i}/\|\hat{f}_{i}\|_{1}\|_{\infty}\leqslant O(\|f_{i}-\hat{f}_{i}\|_{\infty})\leqslant O(L_{0}^{-1/d_{\max}^{\text{in}}}).$
@@
By Theorem 3, there exists a neural network $\hat{g}_{j}$ with architecture in Theorem 3 such that
$W(P(U_{j}),P(\hat{g}_{j}(Z_{j})))\leqslant O\left(L_{1}^{-2/d_{j}^{U}}+L_{2}^{-2/d_{j}^{U}}+(\tau-\tau\log\tau)\right),$
@@
where $Z_{j}\sim U([0,1]^{N_{C,j}})$ i.i.d., $N_{C,j}$ is the number of connected components of support of $U_{j}$, $d_{j}^{U}$ is the dimension of latent variables $U_{j}$ and $L_{1},L_{2},\tau$ are hyperparameters of the neural network defined in Theorem 3. Let $\hat{U}_{j}=\hat{g}_{j}(Z_{j})$, By Lemma 5,
$W(P(U_{1},\cdots,U_{n_{U}}),P(\hat{U}_{1},\cdots,\hat{U}_{n_{U}}))\leqslant\sum^{n_{U}}_{j=1}W(P(U_{j}),P(\hat{g}_{j}(Z_{j})))\leqslant O\left(L_{1}^{-2/d_{j}^{U}}+L_{2}^{-2/d_{j}^{U}}+(\tau-\tau\log\tau)\right).$
@@
Let $\hat{\mathcal{M}}$ be the casual model with the following structure equations
@@
$\hat{V}_{i}=\begin{cases}\hat{f}_{i}\left(\text{Pa}(\hat{V}_{i}),(\hat{g}_{j}(Z_{C_{j}}))_{U_{C_{j}}\in\boldsymbol{U}_{V_{i}}}\right),&V_{i}\text{ is continuous},\\ \arg\max_{k\in[n_{i}]}\left\{g_{k}+\log\left(\hat{f}_{i}\left(\text{Pa}(\hat{V}_{i}),(\hat{g}_{j}(Z_{C_{j}}))_{U_{C_{j}}\in\boldsymbol{U}_{V_{i}}}\right)\right)_{k}\right\},&V_{i}\text{ is categorical}.\end{cases}$
By Theorem 1, we get
@@
$W\left(P^{\mathcal{M}^{*}}(\boldsymbol{V}(\boldsymbol{t})),P^{\hat{\mathcal{M}}}(\boldsymbol{V}(\boldsymbol{t}))\right)$ $\leqslant O(\sum_{V_{i}\ \text{continuous}}\|f_{i}-\hat{f}_{i}\|_{\infty}+\sum_{V_{i}\ \text{categorical}}\|f_{i}-\hat{f}_{i}/\|\hat{f}_{i}\|_{1}\|_{\infty}$
$\quad\quad+W(P(U_{1},\cdots,U_{n_{U}}),P(\hat{U}_{1},\cdots,\hat{U}_{n_{U}})))$

### v2

@@ B.1 Illustration of how to construct canonical representations and neural architectures
$(\gamma_{n+1})_{\#}\lambda(C_{O_{n+1}(i)})=\mathbb{P}(C_{O_{n+1}(i)}),(\gamma_{n+1})_{\#}\lambda(\partial C_{O_{n+1}(i)})=0,\forall i.$ (B.13)
The idea is to construct a piecewise linear curve going through all cubes in $\mathcal{C}_{n+1}$ exactly once and modify its speed. Take $v_{0}\in\text{Center}(C_{O_{n+1}(1)}),v_{i}\in\text{Center}(C_{O_{n+1}(i)}\cap C_{O_{n+1}(i+1)}),v_{2^{d(n+1)}}\in\text{Center}(C_{O_{n+1}(2^{d(n+1)})})$, where $\text{Center}(\prod_{i=1}^{k}[a_{i},b_{i}])=((a_{i}+b_{i})/2)_{i=1,\cdots,k}$ is the center of a cube. $v_{i}$ is well-defined since $C_{O_{n+1}(i)}\cap C_{O_{n+1}(i+1)}$ are cubes of dimension $d-1$ by Lemma B.2. One possible choice of $\gamma_{n+1}(t)$ is
$\gamma_{n+1}(t)=v_{0}+\sum_{i=1}^{2^{(n+1)d}}\frac{1}{p_{i}}(v_{i}-v_{i-1})\left(\left(t-\sum_{k=1}^{i-1}p_{k}\right)^{+}-\left(t-\sum_{k=1}^{i}p_{k}\right)^{+}\right).$ (B.14)
@@
Finally, for any $k_{1}\geqslant k_{2}\geqslant n,t\in[0,1]$, by property 2 in step 1, (B.12) and (B.14), $\gamma_{k_{1}}(t),\gamma_{k_{2}}(t)$ are in one cube $C_{i_{1},i_{2}\cdots,i_{n}}$. and $\text{diam}(C_{i_{1},i_{2}\cdots,i_{n}})\leqslant\left(\frac{2}{3}\right)^{n}\sqrt{d}$. Thus, $|\gamma_{k_{1}}(t)-\gamma_{k_{2}}(t)|\leqslant\left(\frac{2}{3}\right)^{n}\sqrt{d}$ which implies $\{\gamma_{n}(t)\}$ converges uniformly to one continuous function $\gamma(t)$.
@@
$=\sum_{i_{n+1},\cdots,i_{k}}\mathbb{P}(C_{i_{1},i_{2}\cdots,i_{k}})=\mathbb{P}(C_{i_{1},i_{2}\cdots,i_{n}}),$ (B.15)
@@
where we use (B.15) and the fact that $(\gamma_{k})_{\#}\lambda$ vanishes on $\partial C_{i_{1},i_{2}\cdots,i_{k}}$ in the first equality and $\text{diam}(C_{i_{1},i_{2}\cdots,i_{k_{1}}})\leqslant\left(\frac{2}{3}\right)^{k_{1}}\sqrt{d}$ in the last inequality. Let $k\rightarrow\infty$, by Portmanteau Theorem, we get
@@
and we conclude that $\gamma_{\#}\lambda(\partial C_{i_{1},i_{2}\cdots,i_{n}})=0$. By Portmanteau Theorem, let $k\rightarrow\infty$ in (B.15), we obtain $\gamma_{\#}\lambda(C_{i_{1},i_{2}\cdots,i_{n}})=\mathbb{P}(C_{i_{1},i_{2}\cdots,i_{n}})$. Let $\mathcal{C}^{*}=\bigcup_{i=1}^{\infty}\mathcal{C}_{i}$. Notice that for any open set $U\in[0,1]^{d}$, $U=\bigcup_{C\subset U,C\in\mathcal{C}^{*}}C$, which means the $\sigma$-algebra generated by $\mathcal{C}^{*}$ contains all Borel sets. Besides, $\mathcal{C}^{*}$ is a $\pi$-system. Hence, $\lambda\left(\gamma^{-1}(U)\right)=\mathbb{P}(U)$ for all Lebesgue measurable set $U$.
Step 4: Holder Continuity of $\gamma$. We verify the condition of [47, Theorem 1]. $\mathcal{I}_{n}$ and $\mathcal{C}_{n}$ in the above construction correspond to developments $\alpha_{k},\beta_{k}$ in [47, Theorem 1]. Define $f_{k}(I_{O_{k}(i)})=C_{O_{k}(i)},\ i=1,\cdots,2^{kd}$ as a map from $\mathcal{I}_{n}$ to $\mathcal{C}_{n}$.
@@
    If $I_{1},I_{2}\in\mathcal{I}_{k}$ and $I_{1}\cap I_{2}\neq\emptyset$, $I_{1}$, $I_{2}$ are adjacent. By Lemma B.2, $f_{k}(I_{1}),f_{k}(I_{2})$ share a $(d-1)$-dimensional boundary. Therefore, $f_{k}(I_{1})\cap f_{k}(I_{2})\neq\emptyset$.
@@
    Let $A=\prod_{k=1}^{\infty}\left(1+2^{-k}\right)$. We have for any $C\in\mathcal{C}_{n}$, by 3.6
@@
Proof of Theorem 3.8.
The proof is similar to the proof of Theorem 3.5. The difference is that in the first part, we use a deep ReLU network to approximate the Holder continuous curve $\gamma$ instead of wide ReLU network.
Let $\mu_{i}=\mathbb{P}_{C_{i}}/\mathbb{P}(C_{i})$. By 3.4 and 3.6, for each connected component $C_{i}$ of the support, there exists Lipschitz maps $g^{i}_{2}\in\mathcal{H}([0,1]^{d^{C}_{i}},C_{i})\cap\mathcal{F}_{L}([0,1]^{d^{C}_{i}},C_{i})$ such that
@@
for any measurable set $B\subset[0,1]^{d_{i}^{C}}$. By Proposition 3.7, there exist a $1/d_{i}^{C}$-Holder continuous curve $\gamma_{i}:[0,1]\rightarrow[0,1]^{d_{i}^{C}}$ such that $(\gamma_{i})_{\#}\lambda=(g^{i}_{2})^{-1}_{\#}\mu_{i}$.
By [53, Theorem 2] (take $\omega(x)=C\|x\|^{1/d}$ where $C$ is the Holder continuity constant factor), there exists deep ReLu network $\widehat{g}_{1}^{i,j}$ of width $12$ and depth $L_{1}$ such that
$\|\widehat{g}_{1}^{i,j}-(\gamma_{i})_{j}\|_{\infty}\leqslant O\left(L_{1}^{-2/d_{i}^{C}}\right),$
where $(\gamma_{i})_{j}$ is the $j$-coordinate of $\gamma_{i}$. Let $\widehat{g}_{1}^{i}(x)=\left(\widehat{g}_{1}^{i,1}(x),\cdots,\widehat{g}_{1}^{i,d}(x)\right)$, we get
$\|\widehat{g}_{1}^{i}-\gamma_{i}\|_{\infty}\leqslant O\left(L_{1}^{-2/d_{i}^{C}}\right).$
Thus, the width of $\widehat{g}_{1}^{i}$ is $\Theta\left(d_{i}^{C}\right)$. By Lemma E.1
$W\left(\left(g_{2}^{i}\right)^{-1}_{\#}\mu_{i},P(\widehat{g}_{1}^{i}(U_{i}))\right)$ $=W\left((\gamma_{i})_{\#}U_{i},P(\widehat{g}_{1}^{i}(U_{i}))\right)\leqslant\|\widehat{g}_{1}^{i}-\gamma_{i}\|_{\infty}\leqslant O\left(L_{1}^{-2/d_{i}^{C}}\right).$
The rest are the same as proof of Theorem 3.5. ∎
Proof of Corollary 3.9.
@@
$V_{i}=\begin{cases}f_{i}\left(\text{Pa}(V_{i}),\bm{U}_{V_{i}}\right),&V_{i}\ \text{is continuous},\\ \arg\max_{k\in[n_{i}]}\left\{g_{k}^{V_{i}}+\log\left(f_{i}\left(\text{Pa}(V_{i}),\bm{U}_{V_{i}}\right)\right)_{k}\right\},&V_{i}\ \text{is categorical},\|f_{i}\|_{1}=1\end{cases},\quad f_{i}\in\mathcal{F}_{L}.$
@@
$\widehat{f}_{i}=\begin{cases}\arg\min_{\widehat{f}_{i}\in\mathcal{NN}_{d_{i}^{\text{in}},d_{i}^{\text{out}}}(d_{i}^{\text{out}}(2d_{i}^{\text{in}}+10),L_{0})}\|\widehat{f}_{i}-f_{i}\|_{\infty},&V_{i}\ \text{continuous},\\ \max\{0,\arg\min_{\widehat{f}_{i}\in\mathcal{NN}_{d_{i}^{\text{in}},d_{i}^{\text{out}}}(d_{i}^{\text{out}}(2d_{i}^{\text{in}}+10),L_{0})}\|\widehat{f}_{i}-f_{i}\|_{\infty}\},&V_{i}\ \text{categorical}.\end{cases}$
@@
We truncate the neural network at 00 for categorical variables because the propensity functions are required to be non-negative. This truncate operation will not influence the approximation error since $f_{i}$ are non-negative. According to 3.2, $f_{i}$ are Lipschitz continuous. By [53, Theorem 2], we have, if $V_{i}$ is continuous, $\|\widehat{f}_{i}-f_{i}\|_{\infty}\leqslant O(L_{0}^{-1/d_{i}^{\text{in}}})\leqslant O(L_{0}^{-1/d_{\max}^{\text{in}}})$.
If $V_{i}$ is categorical, let $S=\|\widehat{f}\left(\text{pa}(v_{k}),\bm{u}_{V_{k}}\right)-f\left(\text{pa}(v_{k}),\bm{u}_{V_{k}}\right)\|_{\infty}$, $\bm{p}={f}\left(\text{pa}(v_{k}),\bm{u}_{V_{k}}\right),\widehat{\bm{p}}=\widehat{f}\left(\text{pa}(v_{k}),\bm{u}_{V_{k}}\right)/\|\widehat{f}\|_{1}$. If $S>1/(2K)$, $\|\bm{p}-\widehat{\bm{p}}\|_{\infty}\leqslant 2KS$. If $S\leqslant 1/(2K)$,
$\|\bm{p}-\widehat{\bm{p}}\|_{\infty}$ $\leqslant\frac{1}{\|\widehat{f}\|_{1}}\|\widehat{f}\left(\text{pa}(v_{k}),\bm{u}_{V_{k}}\right)-f\left(\text{pa}(v_{k}),\bm{u}_{V_{k}}\right)\|_{\infty}+\frac{1}{\|\widehat{f}\|_{1}}|\|\widehat{f}\|_{1}-1|\|f\|_{\infty}$
$\leqslant\frac{(n_{i}+1)}{\|\widehat{f}\|_{1}}S\leqslant\frac{(n_{i}+1)}{1-n_{i}S}S\leqslant 2(K+1)S$
where we use 3.2 that $n_{i}\leqslant K$ and $S\leqslant 1/(2K)$. Thus,
$\|f_{i}-\widehat{f}_{i}/\|\widehat{f}_{i}\|_{1}\|_{\infty}\leqslant O(\|f_{i}-\widehat{f}_{i}\|_{\infty})\leqslant O(L_{0}^{-1/d_{\max}^{\text{in}}}).$
@@
By Theorem 3.8, there exists a neural network $\widehat{g}_{j}$ with architecture in Theorem 3.8 such that
$W(P(U_{j}),P(\widehat{g}_{j}(Z_{j})))\leqslant O\left(L_{1}^{-2/d_{j}^{U}}+L_{2}^{-2/d_{j}^{U}}+(\tau-\tau\log\tau)\right),$
@@
where $Z_{j}\sim U([0,1]^{N_{C,j}})$ i.i.d., $N_{C,j}$ is the number of connected components of support of $U_{j}$, $d_{j}^{U}$ is the dimension of latent variables $U_{j}$ and $L_{1},L_{2},\tau$ are hyperparameters of the neural network defined in Theorem 3.8. Let $\widehat{U}_{j}=\widehat{g}_{j}(Z_{j})$, By Lemma E.3,
$W(P(U_{1},\cdots,U_{n_{U}}),P(\widehat{U}_{1},\cdots,\widehat{U}_{n_{U}}))\leqslant\sum^{n_{U}}_{j=1}W(P(U_{j}),P(\widehat{g}_{j}(Z_{j})))\leqslant O\left(L_{1}^{-2/d_{j}^{U}}+L_{2}^{-2/d_{j}^{U}}+(\tau-\tau\log\tau)\right).$
@@
Let $\widehat{\mathcal{M}}$ be the casual model with the following structure equations
@@
$\widehat{V}_{i}=\begin{cases}\widehat{f}_{i}\left(\text{Pa}(\widehat{V}_{i}),(\widehat{g}_{j}(Z_{C_{j}}))_{U_{C_{j}}\in\bm{U}_{V_{i}}}\right),&V_{i}\text{ is continuous},\\ \arg\max_{k\in[n_{i}]}\left\{g_{k}+\log\left(\widehat{f}_{i}\left(\text{Pa}(\widehat{V}_{i}),(\widehat{g}_{j}(Z_{C_{j}}))_{U_{C_{j}}\in\bm{U}_{V_{i}}}\right)\right)_{k}\right\},&V_{i}\text{ is categorical}.\end{cases}$
By Theorem 3.3, we get
@@
$W\left(P^{\mathcal{M}^{*}}(\bm{V}(\bm{t})),P^{\widehat{\mathcal{M}}}(\bm{V}(\bm{t}))\right)$ $\leqslant O(\sum_{V_{i}\ \text{continuous}}\|f_{i}-\widehat{f}_{i}\|_{\infty}+\sum_{V_{i}\ \text{categorical}}\|f_{i}-\widehat{f}_{i}/\|\widehat{f}_{i}\|_{1}\|_{\infty}$
$\quad\quad+W(P(U_{1},\cdots,U_{n_{U}}),P(\widehat{U}_{1},\cdots,\widehat{U}_{n_{U}})))$